# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v1

@@ Abstract
Edge-cloud collaborative computing (ECCC) has emerged as a pivotal paradigm for addressing the computational demands of modern intelligent applications, integrating cloud resources with edge devices to enable efficient, low-latency processing. Recent advancements in AI, particularly deep learning and large language models (LLMs), have dramatically enhanced the capabilities of these distributed systems, yet introduce significant challenges in model deployment and resource management. In this survey, we comprehensive examine the intersection of distributed intelligence and model optimization within edge-cloud environments, providing a structured tutorial on fundamental architectures, enabling technologies, and emerging applications. Additionally, we systematically analyze model optimization approaches, including compression, adaptation, and neural architecture search, alongside AI-driven resource management strategies that balance performance, energy efficiency, and latency requirements. We further explore critical aspects of privacy protection and security enhancement within ECCC systems and examines practical deployments through diverse applications, spanning autonomous driving, healthcare, and industrial automation. Performance analysis and benchmarking techniques are also thoroughly explored to establish evaluation standards for these complex systems. Furthermore, the review identifies critical research directions including LLMs deployment, 6G integration, neuromorphic computing, and quantum computing, offering a roadmap for addressing persistent challenges in heterogeneity management, real-time processing, and scalability. By bridging theoretical advancements and practical deployments, this survey offers researchers and practitioners a holistic perspective on leveraging AI to optimize distributed computing environments, fostering innovation in next-generation intelligent systems.
@@ I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, yet with varying scopes and technical emphases. Early works like Abbas et al. [20] and Ferrer et al. [21] establish foundational understanding of edge computing and decentralized cloud architectures respectively, but lack emphasis on AI-driven model optimization. Similarly, Liu et al. [22] provide valuable insights into edge computing systems and tools, while Wang et al. [23] contribute significantly to offloading algorithms, yet neither fully address the integration of AI techniques with edge-cloud architectures. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [24] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management but limited attention to model optimization techniques. Domain-specific surveys like Arthurs et al. [25] focus on intelligent transportation systems without broader applicability across diverse sectors. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain underexplored, with Qu et al. [26] providing a dedicated investigation into mobile edge intelligence for LLMs, yet primarily focusing on resource-efficient techniques rather than holistic system integration across the computing continuum. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications but lack comprehensive treatment of model optimization frameworks. Moreover, while Yang et al. [4] extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, their work predominantly emphasizes network dynamics rather than systematically addressing privacy and cross-domain applications. Even the most recent comprehensive works by Yao et al. [11], Gu et al. [27], and Wang et al. [2] exhibit limitations—addressing either model optimization or edge-cloud collaboration but not systematically integrating with comprehensive privacy frameworks and cross-domain applications. As summarized in Table I, these surveys collectively fail to address the full spectrum of ECCC, particularly in integrating model optimization, privacy, security, and diverse applications. Consequently, our survey stands out by providing a holistic analysis that encompasses the entire computing continuum, advanced technical focuses including LLMs, and practical applications across manufacturing, autonomous driving, and healthcare, thereby filling a critical gap in the literature.
@@ I-D Research Methodology and Organization
Recent analysis reveals significant research momentum in ECCC, as illustrated in Fig. 2, where publication trends and citation patterns demonstrate the evolving landscape across four interrelated domains (The statistics data is collected as of 2025/03/05). While traditional edge computing maintains consistent publication volumes with steadily increasing citation impact, cloud computing exhibits characteristics of a mature research area with decreasing publication numbers yet sustained high citation rates. Most notably, edge-cloud computing shows clear upward trajectories in both publications and citations, signaling growing research interest, whereas the specialized subfield of edge-cloud collaboration displays exponential citation growth peaking in 2024, indicating its emerging importance. Our systematic review methodology employed comprehensive literature searches across prominent scholarly databases including IEEE Xplore, ACM Digital Library, ScienceDirect, and Web of Science, utilizing carefully selected keywords such as ”edge-cloud computing”, ”distributed intelligence”, ”model optimization”, ”AI”, and ”machine learning”. Following retrieval, selected papers pertinent to AI-driven ECCC were carefully reviewed and categorized based on their primary contributions aligned with the scope of this work.
@@ II-C Resource Types and Constraints
Figure 3: Concepts of architecture. The diagram provides a holistic view of the technical scope covered in this survey, where the architecture integrates centralized cloud resources, decentralized edge nodes, and diverse mobile/edge devices. AI-enabled technologies include model compression, adaptation, and resource-aware optimization, which components jointly support distributed intelligence and efficient model execution across multiple application scenarios.
@@ III-A Model Compression
TSCF [51] 2024 IEEE IoTJ P Edge-Cloud Systems Inference Latency $\checkmark$ $\checkmark$
Hybrid SD [52] 2024 ArXiv P Image Synthesis Parameter Efficiency, Image Quality $\checkmark$ $\checkmark$
PSFed [53] 2024 IEEE JSAC P Satellite Edge Cloud Latency, Energy, Privacy $\checkmark$ $\checkmark$
3CE2C [54] 2024 ASC Q Healthcare/ECG Accuracy, Energy Efficiency $\checkmark$ $\checkmark$
CDC [55] 2020 IJCAI P Edge-Cloud Systems Bandwidth Efficiency $\checkmark$ $\checkmark$
QuAsyncFL [56] 2024 IEEE IoTJ Q AIoT Communication Efficiency $\checkmark$ $\checkmark$
Zhang et al. [57] 2025 IEEE JSAC Q Aerial Edge Computing Mean Average Precision (mAP) $\checkmark$ $\checkmark$
DTS-KD [49] 2024 IEEE IoTJ K Industrial IoT F1 Score, Accuracy, Precision, Recall $\checkmark$ $\checkmark$
Wang et al. [58] 2024 CVPR K Multimodal LLMs Generalization Performance $\checkmark$ $\checkmark$
@@
FedAgg [60] 2024 INFOCOM K End-Edge-Cloud Computing Model Accuracy, Convergence Rate $\checkmark$ $\checkmark$
Yang et al. [6] 2025 ADHOC K IoT Traffic Identification Classification Accuracy, Resource Efficiency $\checkmark$ $\checkmark$
GKT [61] 2024 ACL K LLM Inference Inference Speed, Accuracy, Cost $\checkmark$ $\checkmark$
OTO [62] 2024 IEEE TPDS K Reinforcement Learning Convergence Speed, Reward $\checkmark$ $\checkmark$
Zhang et al. [63] 2024 ESWA K Manufacturing Processes Fault Detection Accuracy $\checkmark$ $\checkmark$
@@ III-A1 Pruning
Strategically eliminating redundant model parameters, pruning techniques reduce size and complexity in ECCC, enabling efficient AI model deployment on resource-constrained edge devices [52]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [51], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy and computational cost. For instance, Hybrid SD [52] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [53] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.
@@ III-B Model Adaptation
CoLLaRS [64] 2024 FGCS CL AIoT Applications $\checkmark$ $\checkmark$ $\checkmark$
Li et al. [65] 2024 IEEE TNSM TL Manufacturing $\checkmark$ $\checkmark$ $\checkmark$
Pan et al. [66] 2024 IEEE Netw. FL Large Model Services $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MADDPG [67] 2024 IEEE IoTJ TL 6G Networks $\checkmark$ $\checkmark$ $\checkmark$
DTS-KD [49] 2024 IEEE IoTJ TL Industrial IoT $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Wang et al. [58] 2024 CVPR CL Multimodal Large Language Models $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Wang et al. [68] 2024 ESWA FL Rotating Machinery Fault Diagnosis $\checkmark$ $\checkmark$ $\checkmark$
FedAgg [60] 2024 INFOCOM FL End-Edge-Cloud Systems $\checkmark$ $\checkmark$ $\checkmark$
OTO [62] 2024 IEEE TPDS TL Cloud-Edge Collaborative DRL $\checkmark$ $\checkmark$ $\checkmark$
U-Federator [69] 2024 JNCA FL Cloud-Edge-Fog Computing $\checkmark$ $\checkmark$ $\checkmark$
OTFAC [70] 2024 IEEE IoTJ FL Cloud-Edge Collaborative IoT $\checkmark$ $\checkmark$ $\checkmark$
FLAT [71] 2024 IEEE ISCC FL Cloud-Edge-Client Continuum $\checkmark$ $\checkmark$ $\checkmark$
FedITA [72] 2024 AEI FL Industrial Motor Fault Diagnosis $\checkmark$ $\checkmark$ $\checkmark$
Khuen et al. [73] 2024 IEEE Access FL Precision Aquaculture $\checkmark$ $\checkmark$ $\checkmark$
Kim et al. [74] 2024 IEEE IoTJ FL Cloud-Edge-Terminal IoT Networks $\checkmark$ $\checkmark$ $\checkmark$
QuAsyncFL [56] 2024 IEEE IoTJ FL Cloud-Edge-Terminal AIoT $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Loconte et al. [75] 2024 FGCS FL Cloud-to-Edge-IoT Continuum $\checkmark$ $\checkmark$ $\checkmark$
BAFL-DT [76] 2024 IEEE TSC FL Digital Twin Edge-Cloud Continuum $\checkmark$ $\checkmark$ $\checkmark$
FDRL [77] 2024 IJPR FL Cloud-Edge Manufacturing Systems $\checkmark$ $\checkmark$ $\checkmark$
Wei et al. [78] 2024 IEEE TPDS FL Edge Clouds $\checkmark$ $\checkmark$ $\checkmark$
Yang et al. [79] 2024 IEEE IoTJ FL AIoT Data Prediction $\checkmark$ $\checkmark$ $\checkmark$
FedCAE [80] 2024 IEEE TIE FL Machine Fault Diagnosis $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FlexibleFL [81] 2024 INS FL Cloud-Edge Federated Learning $\checkmark$ $\checkmark$ $\checkmark$
PSPFL [82] 2024 IEEE TMC FL Mobile Edge-Cloud Networks $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CCSFLF [83] 2024 CCPE FL Federated Learning Systems $\checkmark$ $\checkmark$ $\checkmark$
Zhou et al. [13] 2024 ADHOC FL AIoT Quality of Service $\checkmark$ $\checkmark$ $\checkmark$
CADCL-DTM [84] 2024 CAC CL Rolling Bearings Monitoring $\checkmark$ $\checkmark$
Li et al. [85] 2024 IEEE TMC CL Edge Computing Service Provisioning $\checkmark$ $\checkmark$ $\checkmark$
Lian et al. [86] 2024 SDI CL ITS Object Detection $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
EdgeC3 [87] 2024 IEEE TMC CL Distributed Deep Learning $\checkmark$ $\checkmark$ $\checkmark$
@@ III-C3 Reinforcement Learning for Model Optimization
In dynamic edge-cloud systems, RL provides a powerful mechanism for optimizing model parameters and offloading decisions [101]. RL agents learn optimal policies by leveraging environmental feedback, adapting model configurations and resource allocation strategies to improve system adaptability and performance. Specifically, techniques like deep Q-learning [102] and asynchronous proximal policy Optimization (APPO) [103] guide agents to explore actions and learn strategies that maximize cumulative rewards. For example, a study in [18] combined deep reinforcement learning with edge computing to optimize real-time monitoring and control in IoT environments, ultimately improving system stability and reducing failure rates. Additionally, RL enables sophisticated offloading mechanisms that consider factors such as latency, energy consumption, and task dependencies [104]. Through dynamic parameter adjustment, such as the number of layers or computational precision, RL optimizes the trade-off between model accuracy and computational cost [105]. Similarly, RL-based task offloading strategies dynamically determine optimal execution locations (edge or cloud) by adapting to real-time network and system conditions [103], simultaneously maximizing resource efficiency and reducing latency for time-critical applications.
@@ IV-A Task Offloading Mechanisms
Zhang et al. [19] Edge Cloud Architecture Time, Cost, Load balance, User satisfaction, Trust measurement Found more than half of the best solutions on benchmark functions $\checkmark$ $\checkmark$
@@
Fan et al. [118] Edge-Cloud Cooperative Computing Task processing delay, Task queuing stability Superior performance compared to 4 other schemes $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Qi et al. [92] Edge-Cloud Service Provisioning Service quality, Cost optimization Effective cost-service quality balance demonstrated $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Abouaomar et al. [119] Edge Computing for IoT Latency, Resource consumption Outperforms benchmark approaches in latency and resource utilization $\checkmark$ $\checkmark$
@@
Li et al. [120] Edge-Cloud Cooperation Long-term network performance, Energy constraints Converges to approximate optimal solution in polynomial time $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Luckow et al. [121] Edge-to-Cloud Applications Performance impact of task placement 5-65% performance improvement depending on scenario $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Cai et al. [122] Data-Intensive Edge Services Throughput, Delay, Resource consumption Superior performance compared to state-of-the-art algorithms $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Pang et al. [123] Vehicle-Edge-Cloud Latency, Energy consumption Significant decrease in latency and energy consumption $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
PRAISE [124] Multi-Access Edge Cloud System benefits, Resource costs Outperforms baselines with higher benefits and lower costs $\checkmark$ $\checkmark$ $\checkmark$
@@
Khalafi et al. [125] Edge-Cloud Orchestrated Networks System cost, QoS, Delay violations Reduced costs associated with network slicing $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Peng et al. [126] Edge Computing Systems Job computing time, Job blocking probability Optimal replication factor varies with distribution parameters and arrival rate $\checkmark$ $\checkmark$ $\checkmark$
@@
DRLIS [127] Edge-Fog-Cloud Computing Load balancing, Response time, Weighted cost Reduces execution cost by up to 55% (load balancing), 37% (response time), 50% (weighted cost) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Hua et al. [96] Heterogeneous Edge-Cloud Computing Energy consumption Significantly reduced energy consumption of mobile devices $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Arroba et al. [128] Sustainable Edge Computing Energy efficiency, Economic viability Early results demonstrate benefits for future computing environments $\checkmark$ $\checkmark$
@@
PDCO [129] Cloud-Edge Collaboration Total energy consumption, Resource utilization Approaches optimal performance in energy consumption and resource utilization $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
He et al. [130] Wireless Powered MEC Systems Amount of processed data Enhances computation performance compared to benchmark methods $\checkmark$ $\checkmark$
@@
iGATS [8] IoT-Edge-Cloud Networks Response latency, Temporal satisfaction, Load-balancing Outperforms baseline techniques for reducing response latency $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
ETCRA [94] Edge-Cloud Workflows Comprehensive system function, Execution time, Energy consumption, Reliability Significantly outperforms other baselines on four key measurements $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Hu et al. [131] MEC-enabled Mobile IoT Energy efficiency, Service delay Achieves [O(1/V), O(V)] tradeoff between EE and service delay $\checkmark$ $\checkmark$
@@
Alopex-DE [111] Local-Edge-Cloud Computing Latency, Energy consumption Superior effectiveness compared to three state-of-the-art algorithms $\checkmark$ $\checkmark$ $\checkmark$
@@
DT-MADDPG [97] Digital Twin Enabled 6G IIoT Task success rate, Latency, Energy consumption Significantly improved task success rate with reduced latency and energy consumption $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Fan et al. [132] Cloud-Edge-End Cooperative Networks Task processing delay, Energy consumption, Block mining speed Superior performance compared to 5 reference schemes in 7 scenarios $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Zhang et al. [133] Edge-Cloud Systems Task completion time, Cost, Energy consumption, System reliability Outperforms existing design approaches $\checkmark$ $\checkmark$ $\checkmark$
@@
Qi et al. [92] Edge-Cloud Collaboration Cost, Service quality Effective in both synthetic and real-world data experiments $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
EdgeOptimizer [134] Kubernetes-Based Edge-Cloud Clusters Success rate of time-critical tasks, Overall overload Improved success rate for time-critical tasks $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$

### v2

@@ Abstract
Edge-cloud collaborative computing (ECCC) has emerged as a pivotal paradigm for addressing the computational demands of modern intelligent applications, integrating cloud resources with edge devices to enable efficient, low-latency processing across distributed communication networks. Recent advancements in AI, particularly deep learning and large language models (LLMs), have dramatically enhanced the capabilities of these networked systems, yet introduce significant challenges in model deployment, network resource management, and cross-layer optimization. In this survey, we comprehensively examine the intersection of distributed intelligence and model optimization within edge-cloud environments, providing a structured tutorial on fundamental architectures, communication protocols, and network-aware computing frameworks. Additionally, we systematically analyze model optimization approaches, including compression, adaptation, and neural architecture search, alongside AI-driven resource management strategies that balance performance, energy efficiency, and communication overhead across heterogeneous networks. We further explore critical aspects of privacy protection and security enhancement within ECCC systems and examine practical deployments through diverse networked applications, spanning autonomous driving, healthcare, and industrial automation. Performance analysis and benchmarking techniques are also thoroughly explored to establish evaluation standards for these complex distributed systems. Furthermore, the review identifies critical research directions including LLMs deployment, 6G integration, neuromorphic computing, and quantum computing, offering a roadmap for addressing persistent challenges in heterogeneity management, real-time processing, and scalability. By bridging theoretical advancements in communications with practical deployments, this survey offers researchers and practitioners a holistic perspective on leveraging AI to optimize distributed computing environments over next-generation communication networks, fostering innovation in intelligent networked systems.
@@ I-A Background
The evolution of communication networks plays a pivotal role in enabling effective edge-cloud collaboration by providing the foundational infrastructure that connects distributed computational resources. Recent advancements in network technologies, particularly software-defined networking (SDN) [20] and network function virtualization (NFV) [21], have significantly enhanced the flexibility and programmability of communication infrastructures supporting ECCC. These technologies facilitate dynamic network resource allocation and traffic engineering, which are essential for optimizing data flows between edge devices and cloud servers [22]. Furthermore, emerging wireless communication paradigms, including millimeter-wave communications, massive MIMO, and non-orthogonal multiple access (NOMA) [23], are revolutionizing the capacity, reliability, and efficiency of edge-cloud connectivity. The integration of network slicing techniques [24] enables the creation of virtualized, isolated network segments with tailored quality-of-service guarantees for diverse edge-cloud applications. Moreover, cross-layer optimization approaches that jointly consider communication, computation, and caching resources [25] are becoming increasingly important for minimizing end-to-end latency and maximizing system throughput. As communication networks evolve toward 6G, the synergistic optimization of networking and computing resources will be instrumental in addressing the growing demands of distributed intelligence across the edge-cloud continuum [26].
@@ I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, yet with varying scopes and technical emphases. Early works like Abbas et al. [27] and Ferrer et al. [28] establish a foundational understanding of edge computing and decentralized cloud architectures respectively, but lack emphasis on AI-driven model optimization. Similarly, Liu et al. [29] provide valuable insights into edge computing systems and tools, while Wang et al. [30] contribute significantly to offloading algorithms, yet neither fully addresses the integration of AI techniques with edge-cloud architectures. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [31] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management but limited attention to model optimization techniques. Domain-specific surveys like Arthurs et al. [32] focus on intelligent transportation systems without broader applicability across diverse sectors. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain underexplored, with Qu et al. [33] providing a dedicated investigation into mobile edge intelligence for LLMs, yet primarily focusing on resource-efficient techniques rather than holistic system integration across the computing continuum. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications but lack comprehensive treatment of model optimization frameworks. Moreover, while Yang et al. [4] extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, their work predominantly emphasizes network dynamics rather than systematically addressing privacy and cross-domain applications. Even the most recent comprehensive works by Yao et al. [11], Gu et al. [34], and Wang et al. [2] exhibit limitations—addressing either model optimization or edge-cloud collaboration but not systematically integrating with comprehensive privacy frameworks and cross-domain applications. As summarized in Table I, these surveys collectively fail to address the full spectrum of ECCC, particularly in integrating model optimization, privacy, security, and diverse applications. Consequently, our survey stands out by providing a holistic analysis that encompasses the entire computing continuum, advanced technical focuses including LLMs, and practical applications across manufacturing, autonomous driving, and healthcare, thereby filling a critical gap in the literature.
@@ I-D Research Methodology and Organization
Recent analysis reveals significant research momentum in ECCC, as illustrated in Fig. 2, where publication trends and citation patterns demonstrate the evolving landscape across four interrelated domains (The statistics data is collected as of 2025/03/05). While traditional edge computing maintains consistent publication volumes with steadily increasing citation impact, cloud computing exhibits characteristics of a mature research area with decreasing publication numbers yet sustained high citation rates. Most notably, edge-cloud computing shows clear upward trajectories in both publications and citations, signaling growing research interest, whereas the specialized subfield of edge-cloud collaboration displays exponential citation growth peaking in 2024, indicating its emerging importance. Our systematic review methodology employed comprehensive literature searches across prominent scholarly databases including IEEE Xplore, ACM Digital Library, ScienceDirect, and Web of Science, utilizing carefully selected keywords such as ”edge-cloud computing”, ”distributed intelligence”, ”model optimization”, ”AI”, and ”machine learning”. Following retrieval, selected papers pertinent to ECCC were carefully reviewed and categorized based on their primary contributions aligned with the scope of this work.
@@ II-C Resource Types and Constraints
Figure 3: Concepts of architecture. The diagram provides a holistic view of the technical scope covered in this survey, where the architecture integrates centralized cloud resources, decentralized edge nodes, and diverse mobile/edge devices. AI-enabled technologies include model compression, adaptation, and resource-aware optimization, whose components jointly support distributed intelligence and efficient model execution across multiple application scenarios.
@@ III-A Model Compression
TSCF [58] 2024 IEEE IoTJ P Edge-Cloud Systems Inference Latency
Hybrid SD [59] 2024 ArXiv P Image Synthesis Parameter Efficiency, Image Quality
PSFed [60] 2024 IEEE JSAC P Satellite Edge Cloud Latency, Energy, Privacy
3CE2C [61] 2024 ASC Q Healthcare/ECG Accuracy, Energy Efficiency
CDC [62] 2020 IJCAI P Edge-Cloud Systems Bandwidth Efficiency
QuAsyncFL [63] 2024 IEEE IoTJ Q AIoT Communication Efficiency
Zhang et al. [64] 2025 IEEE JSAC Q Aerial Edge Computing Mean Average Precision (mAP)
DTS-KD [56] 2024 IEEE IoTJ K Industrial IoT F1 Score, Accuracy, Precision, Recall
Wang et al. [65] 2024 CVPR K Multimodal LLMs Generalization Performance
@@
FedAgg [67] 2024 INFOCOM K End-Edge-Cloud Computing Model Accuracy, Convergence Rate
Yang et al. [6] 2025 ADHOC K IoT Traffic Identification Classification Accuracy, Resource Efficiency
GKT [68] 2024 ACL K LLM Inference Inference Speed, Accuracy, Cost
OTO [69] 2024 IEEE TPDS K Reinforcement Learning Convergence Speed, Reward
Zhang et al. [70] 2024 ESWA K Manufacturing Processes Fault Detection Accuracy
@@ III-A1 Pruning
Strategically eliminating redundant model parameters, pruning techniques reduce size and complexity in ECCC, enabling efficient AI model deployment on resource-constrained edge devices [59]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [58], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy, and computational cost. For instance, Hybrid SD [59] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [60] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.
@@ III-B Model Adaptation
CoLLaRS [71] 2024 FGCS CL AIoT Applications
Li et al. [72] 2024 IEEE TNSM TL Manufacturing
Pan et al. [73] 2024 IEEE Netw. FL Large Model Services
MADDPG [74] 2024 IEEE IoTJ TL 6G Networks
DTS-KD [56] 2024 IEEE IoTJ TL Industrial IoT
Wang et al. [65] 2024 CVPR CL Multimodal Large Language Models
Wang et al. [75] 2024 ESWA FL Rotating Machinery Fault Diagnosis
FedAgg [67] 2024 INFOCOM FL End-Edge-Cloud Systems
OTO [69] 2024 IEEE TPDS TL Cloud-Edge Collaborative DRL
U-Federator [76] 2024 JNCA FL Cloud-Edge-Fog Computing
OTFAC [77] 2024 IEEE IoTJ FL Cloud-Edge Collaborative IoT
FLAT [78] 2024 IEEE ISCC FL Cloud-Edge-Client Continuum
FedITA [79] 2024 AEI FL Industrial Motor Fault Diagnosis
Khuen et al. [80] 2024 IEEE Access FL Precision Aquaculture
Kim et al. [81] 2024 IEEE IoTJ FL Cloud-Edge-Terminal IoT Networks
QuAsyncFL [63] 2024 IEEE IoTJ FL Cloud-Edge-Terminal AIoT
Loconte et al. [82] 2024 FGCS FL Cloud-to-Edge-IoT Continuum
BAFL-DT [83] 2024 IEEE TSC FL Digital Twin Edge-Cloud Continuum
FDRL [84] 2024 IJPR FL Cloud-Edge Manufacturing Systems
Wei et al. [85] 2024 IEEE TPDS FL Edge Clouds
Yang et al. [86] 2024 IEEE IoTJ FL AIoT Data Prediction
FedCAE [87] 2024 IEEE TIE FL Machine Fault Diagnosis
FlexibleFL [88] 2024 INS FL Cloud-Edge Federated Learning
PSPFL [89] 2024 IEEE TMC FL Mobile Edge-Cloud Networks
CCSFLF [90] 2024 CCPE FL Federated Learning Systems
Zhou et al. [13] 2024 ADHOC FL AIoT Quality of Service
CADCL-DTM [91] 2024 CAC CL Rolling Bearings Monitoring
Li et al. [92] 2024 IEEE TMC CL Edge Computing Service Provisioning
Lian et al. [93] 2024 SDI CL ITS Object Detection
EdgeC3 [94] 2024 IEEE TMC CL Distributed Deep Learning
@@ III-C3 Reinforcement Learning for Model Optimization
In dynamic edge-cloud systems, RL provides a powerful mechanism for optimizing model parameters and offloading decisions [107]. RL agents learn optimal policies by leveraging environmental feedback, adapting model configurations and resource allocation strategies to improve system adaptability and performance. Specifically, techniques like deep Q-learning [108] and asynchronous proximal policy Optimization (APPO) [109] guide agents to explore actions and learn strategies that maximize cumulative rewards. For example, a study in [18] combined deep reinforcement learning with edge computing to optimize real-time monitoring and control in IoT environments, ultimately improving system stability and reducing failure rates. Additionally, RL enables sophisticated offloading mechanisms that consider factors such as latency, energy consumption, and task dependencies [110]. Through dynamic parameter adjustments, such as the number of layers or computational precision, RL optimizes the trade-off between model accuracy and computational cost [111]. Similarly, RL-based task offloading strategies dynamically determine optimal execution locations (edge or cloud) by adapting to real-time network and system conditions [109], simultaneously maximizing resource efficiency and reducing latency for time-critical applications.
@@ IV-A Task Offloading Mechanisms
Zhang et al. [19] Edge Cloud Architecture Time, Cost, Load balance, User satisfaction, Trust measurement Found more than half of the best solutions on benchmark functions
@@
Fan et al. [123] Edge-Cloud Cooperative Computing Task processing delay, Task queuing stability Superior performance compared to 4 other schemes
@@
Qi et al. [99] Edge-Cloud Service Provisioning Service quality, Cost optimization Effective cost-service quality balance demonstrated
@@
Abouaomar et al. [124] Edge Computing for IoT Latency, Resource consumption Outperforms benchmark approaches in latency and resource utilization
@@
Li et al. [125] Edge-Cloud Cooperation Long-term network performance, Energy constraints Converges to approximate optimal solution in polynomial time
@@
Luckow et al. [126] Edge-to-Cloud Applications Performance impact of task placement 5-65% performance improvement depending on scenario
@@
Cai et al. [127] Data-Intensive Edge Services Throughput, Delay, Resource consumption Superior performance compared to state-of-the-art algorithms
@@
Pang et al. [128] Vehicle-Edge-Cloud Latency, Energy consumption Significant decrease in latency and energy consumption
@@
PRAISE [129] Multi-Access Edge Cloud System benefits, Resource costs Outperforms baselines with higher benefits and lower costs
@@
Khalafi et al. [130] Edge-Cloud Orchestrated Networks System cost, QoS, Delay violations Reduced costs associated with network slicing
@@
Peng et al. [131] Edge Computing Systems Job computing time, Job blocking probability Optimal replication factor varies with distribution parameters and arrival rate
@@
DRLIS [132] Edge-Fog-Cloud Computing Load balancing, Response time, Weighted cost Reduces execution cost by up to 55% (load balancing), 37% (response time), 50% (weighted cost)
@@
Hua et al. [102] Heterogeneous Edge-Cloud Computing Energy consumption Significantly reduced energy consumption of mobile devices
@@
Arroba et al. [133] Sustainable Edge Computing Energy efficiency, Economic viability Early results demonstrate benefits for future computing environments
@@
PDCO [134] Cloud-Edge Collaboration Total energy consumption, Resource utilization Approaches optimal performance in energy consumption and resource utilization
@@
He et al. [135] Wireless Powered MEC Systems Amount of processed data Enhances computation performance compared to benchmark methods
@@
iGATS [8] IoT-Edge-Cloud Networks Response latency, Temporal satisfaction, Load-balancing Outperforms baseline techniques for reducing response latency
@@
ETCRA [101] Edge-Cloud Workflows Comprehensive system function, Execution time, Energy consumption, Reliability Significantly outperforms other baselines on four key measurements
@@
Hu et al. [136] MEC-enabled Mobile IoT Energy efficiency, Service delay Achieves [O(1/V), O(V)] tradeoff between EE and service delay
@@
Alopex-DE [116] Local-Edge-Cloud Computing Latency, Energy consumption Superior effectiveness compared to three state-of-the-art algorithms
@@
DT-MADDPG [103] Digital Twin Enabled 6G IIoT Task success rate, Latency, Energy consumption Significantly improved task success rate with reduced latency and energy consumption
@@
Fan et al. [137] Cloud-Edge-End Cooperative Networks Task processing delay, Energy consumption, Block mining speed Superior performance compared to 5 reference schemes in 7 scenarios
@@
Zhang et al. [138] Edge-Cloud Systems Task completion time, Cost, Energy consumption, System reliability Outperforms existing design approaches
@@
Qi et al. [99] Edge-Cloud Collaboration Cost, Service quality Effective in both synthetic and real-world data experiments
@@
EdgeOptimizer [139] Kubernetes-Based Edge-Cloud Clusters Success rate of time-critical tasks, Overall overload Improved success rate for time-critical tasks